OpenStudy (anonymous):

I need a refresher. If you have 6(x-1) you get 6x and then do you multiply 6 (-1) and what are the rules again is it -6 or +6?

OpenStudy (accessdenied):

6 times -1 is -6 if we have a negative times a positive, then it is negative if we have a positive times a positive or negative times a negative, then it becomes positive

OpenStudy (anonymous):

+*+=+ -*+=- -*-=+

OpenStudy (accessdenied):

and yes, that would be the correct step -- multiplying 6 and -1 together just wanted to make sure that was clear we're using distributive property, so we multiply the outer number by each term inside the parentheses.

OpenStudy (anonymous):

Thank you!!! what about a + times a -?

OpenStudy (accessdenied):

a positive times a negative results in a negative ie. -1*6 = -6, -2 * 2 = -4, -x * 3 = -3x

OpenStudy (anonymous):

it would still be negative, because you are multiplying and it does not change

OpenStudy (anonymous):

Ok. :)

OpenStudy (accessdenied):

positive times a negative and negative times a positive are the same from multiplication properties a*b = b*a (commutative)